# Universal Adaptive Constraint Propagation: Scaling Structured Inference for Large Language Models via Meta-Reinforcement Learning

### v1

@@ 4.6 Mechanistic Interpretability: Probing Studies
On sentences with nested relative clauses (e.g., “The manager [who hired the employee [who left]] resigned”), the policy discovers a middle-out propagation strategy. Rather than processing constraints bottom-up (inner clauses first) or top-down (outer attachment first), the policy establishes clause boundaries at mid-depth first, reducing backtracks from 2.3 to 1.5 on average compared to standard heuristics. This strategy emerges without explicit supervision, suggesting the policy discovers structural regularities through reward optimization alone. The middle-out strategy processes constraints at depths 2–3 first (relative clause boundaries), establishing structural anchors before resolving attachment ambiguities at depths 1 and 4+, which reduces backtracking by 34% compared to bottom-up processing. A visualization of this depth-ordered propagation pattern is provided in the supplementary materials.
4.7 Linguistic Error Analysis and Cross-Task Transfer
@@
We characterize error cases by sentence length, ambiguity, constraint density, and schema depth. Errors concentrate on sentences with high ambiguity (coordination, attachment) and long-distance dependencies. For transfer stress tests, we evaluate hard transfer directions: TSP to Scheduling achieves only 74% of specialist performance with 5–10 steps, but recovers to 87% with 20–30 gradient steps or a small adapter layer (2-layer MLP, 5K parameters). To test robustness to determinism violations, we simulate stochastic propagators (randomized pruning order with 10% noise) and find that the policy maintains 92% of deterministic performance, with entropy-aware fallback providing additional robustness.
@@ 4.8 Scaling to Long Sentences
5 Generalization Beyond Language: CP Benchmarks
| Solver | Avg. Gap (%) | Runtime (Norm.) | Solve Rate (%) |
| OR-Tools | 8.5 | 1.0 | 88% |
| JuLS (Base) | 9.1 | 1.1 | 84% |
| Activity-Based | 7.8 | 0.95 | 90% |
| MetaJuLS (Ours) | 6.6 | 0.63 | 94% |
5.1 Policy Analysis and Transfer
6 Discussion and Limitations
The success of MetaJuLS demonstrates that reinforcement learning can effectively learn adaptive propagation policies that outperform traditional handcrafted heuristics (silver2016alphago; schulman2017proximal; silver2017alphago). By framing propagation scheduling as a sequential decision problem, where the value of each propagation step depends on the entire search trajectory rather than only on local features (sutton2018reinforcement), we enable the discovery of sophisticated scheduling strategies that adapt to both problem structure and inference state in NLP and in more classical CP settings (van2006propagation; bessiere2006constraint; debruyne1997optimal).
The analysis of learned policies reveals that in simple regimes, the agent rediscovers classical heuristics like Domain-over-Weight (apt2003principles; marriott1998programming), validating that the learning process can recover known effective strategies. In more complex structured problems, the agent discovers novel strategies not present in standard heuristics, such as focusing on violation potentials in graph coloring (regin1994filtering; gent2006watched). The transfer learning results show that policies trained on diverse problem sets generalize to novel domains with minimal performance degradation, which is crucial for practical deployment where problem instances may not match the exact distribution of training data.
At the same time, several limitations should be acknowledged. We evaluate on English parsing and constrained generation, and the approach may not transfer to morphologically rich languages or other structured prediction tasks without adaptation. Meta-training requires 120 GPU-hours on 8 A100 GPUs (one-time cost), enabling rapid adaptation to new tasks with 5–10 gradient steps. For deployment regimes: (1) Low volume (1K queries/day): ROI is negative, but value comes from latency-SLA enablement; (2) High volume (100K queries/day): speedup translates to increased capacity (10K 17K queries/day with same hardware), with ROI break-even at 200 days assuming $2/GPU-hour; (3) With pre-trained checkpoints: zero training cost, immediate deployment. The primary value proposition is latency reduction for user-facing applications: MetaJuLS reduces p99 latency from 245ms to 142ms on constrained generation tasks, enabling deployment in latency-sensitive settings where baseline approaches fail SLA requirements. We provide pre-trained checkpoints enabling deployment without retraining. Our gains are relative to a specific CRF-based chart parser, and modern neural parsers using different inference strategies may not benefit equally. Finally, casting a new NLP task as constraint propagation requires manual effort to define variables, domains, and constraints, though the meta-learning framework reduces the adaptation burden once the constraint formulation is established.
There are also limitations in terms of transfer and robustness. While rapid adaptation achieves 85–92% of specialist performance in many cases, there are domains where transfer performs significantly worse; for example, transferring from TSP to Scheduling achieves only 74% of specialist performance, indicating that some domain-specific knowledge may still be necessary for optimal results. Moreover, the current approach assumes that constraint propagators are deterministic and that their effects can be accurately predicted from the current inference state. In practice, some propagators may have non-deterministic behavior or may interact in complex ways that are difficult to model. Extending the framework to handle uncertainty in propagation effects and to reason about stochastic or approximate propagators is an important direction for future research.
7 Conclusion
We have introduced MetaJuLS, a meta-reinforcement learning framework for adaptive constraint propagation that scales to modern LLM inference. Our contributions are: (1) A unified constraint-propagation scheduling formulation spanning constituency parsing, dependency parsing, and LLM constrained decoding, showing that scheduling order is a first-order efficiency lever in pruned inference regimes; (2) Meta-learning for rapid adaptation across languages and tasks, enabling few-step adaptation (5–10 gradient steps, 5–15 seconds) rather than hours of task-specific training; (3) A safety-aware hybrid policy with principled fallback guarantees, reducing accuracy gaps from 2.1% to 0.15% while maintaining 1.4 average speedups; (4) Ablation-backed evidence that RL is necessary: MetaJuLS outperforms strong heuristic baselines (VSIDS-style, cost-normalized greedy, supervised ranking) by 1.2–1.5 in speed while matching or exceeding accuracy. By combining meta-learning with graph attention networks, MetaJuLS achieves 1.5–2.0 speedups over GPU-optimized baselines while maintaining within 0.2% accuracy of state-of-the-art parsers.
Our mechanistic interpretability analysis reveals that the policy discovers both human-like parsing strategies (easy-first) and novel non-intuitive heuristics (middle-out propagation for nested structures). The safety-aware fallback mechanism ensures accuracy parity while maintaining speedups on 80% of instances. On LLM-constrained generation (LogicBench, GSM8K-Constrained), MetaJuLS outperforms speculative decoding by 1.6–1.8 while improving constraint satisfaction rates, with factorial experiments demonstrating orthogonality when combined.
These results demonstrate that learned constraint propagation can significantly improve efficiency in structured inference tasks. By reducing propagation steps in LLM deployments, MetaJuLS reduces inference latency, enabling deployment in latency-sensitive applications. The meta-learning framework enables rapid adaptation to new tasks with minimal compute, though optimal performance may still require domain-specific knowledge in some cases. Future work could explore extending the framework to other aspects of solver design, investigating the theoretical foundations of rapid adaptation, and developing automated methods for casting new tasks into the constraint-propagation interface.
Reproducibility Statement
Code, trained models, and experiment scripts are available at https://github.com/[anonymous]/metajuls. We include: (1) Full training and evaluation code; (2) Pre-trained meta-learned policy checkpoints; (3) Scripts to reproduce all tables and figures; (4) Preprocessed data splits (PTB, UD, and LLM datasets). Meta-training requires approximately 120 GPU-hours on 8 A100 GPUs (one-time cost), enabling rapid adaptation to new tasks with 5–10 gradient steps (5–15 seconds on a single GPU). Inference adds 3% overhead to base parser/LLM runtime. All experiments use identical hardware (A100 GPUs), batch sizes (1 for parsing, 8 for constrained decoding), and measurement protocols (warmup runs excluded, median of 5 runs).
Impact Statement
MetaJuLS contributes to Green AI by reducing the computational cost of structured inference in LLM deployments. By reducing propagation steps by 1.5–2.0, MetaJuLS directly reduces carbon footprint: for a deployment processing 1M structured queries/day, this translates to approximately 2.4 MWh saved annually under conservative assumptions (100W per query, 50% of queries benefiting from full speedup); actual savings vary with hardware and workload characteristics. The meta-learning framework further reduces environmental impact by eliminating the need for task-specific retraining, enabling a single universal policy to serve multiple languages and domains.
Acknowledgments
We thank the anonymous reviewers for their valuable feedback and suggestions. This work was supported by [funding information]. We also acknowledge the developers of the MiniZinc and XCSP benchmark suites for providing standardized evaluation datasets (van2011minizinc; lecoutre2011xcsp).
Appendix A Implementation Details
@@ A.1 Hyperparameters and Robustness
The reward function hyperparameters are set to and based on preliminary experiments that balanced domain reduction and the deterministic cost proxy described in the main text (sutton2018reinforcement; schulman2017proximal). The discount factor encourages long-term planning while preventing excessive focus on distant future rewards (schulman2017proximal; silver2017alphago).
Appendix B Additional Experimental Results
@@ B.1 NLP Ablation Studies
B.2 Additional Baselines and Ablations
@@
Appendix C Policy Visualization

### v2

@@ Appendix B Mechanistic Analysis and Policy Interpretability
On sentences with nested relative clauses (e.g., “The manager [who hired the employee [who left]] resigned”), the policy discovers a middle-out propagation strategy. Rather than processing constraints bottom-up (inner clauses first) or top-down (outer attachment first), the policy establishes clause boundaries at mid-depth first, reducing backtracks from 2.3 to 1.5 on average compared to standard heuristics. This strategy emerges without explicit supervision, suggesting the policy discovers structural regularities through reward optimization alone shihab2025detecting. The middle-out strategy processes constraints at depths 2–3 first (relative clause boundaries), establishing structural anchors before resolving attachment ambiguities at depths 1 and 4+, which reduces backtracking by 34% compared to bottom-up processing. A visualization of this depth-ordered propagation pattern is provided in the supplementary materials.
B.1 Linguistic Error Analysis
@@
We characterize error cases by sentence length, ambiguity, constraint density, and schema depth. Errors concentrate on sentences with high ambiguity (coordination, attachment) and long-distance dependencies. For transfer stress tests, we evaluate hard transfer directions: TSP to Scheduling achieves only 74% of specialist performance with 5–10 steps, but recovers to 87% with 20–30 gradient steps or a small adapter layer (2-layer MLP, 5K parameters). To test robustness to determinism violations, we simulate stochastic propagators (randomized pruning order with 10% noise) and find that the policy maintains 92% of deterministic performance, with entropy-aware fallback shihab2025differentiable providing additional robustness.
@@ B.2 Scaling Analysis
Appendix C Implementation Details
@@ C.1 Hyperparameters and Robustness
Table 8 details the full model configuration. The reward function hyperparameters are set to and based on preliminary experiments that balanced domain reduction and the deterministic cost proxy described in the main text (sutton2018reinforcement; schulman2017proximal). The discount factor encourages long-term planning while preventing excessive focus on distant future rewards (schulman2017proximal; silver2017alphago).
Appendix D Constraint Programming Experiments: Extended Results
D.1 Policy Analysis and Transfer
Appendix E Additional Experimental Results
@@ E.1 NLP Ablation Studies
E.2 Additional Baselines and Ablations
@@
Appendix F Policy Visualization
@@
Appendix G Deployment Analysis and Transfer Limitations
G.1 Deployment ROI Analysis
For deployment regimes: (1) Low volume (1K queries/day): ROI is negative, but value comes from latency-SLA enablement; (2) High volume (100K queries/day): speedup translates to increased capacity (10K 17K queries/day with same hardware), with ROI break-even at 200 days assuming $2/GPU-hour; (3) With pre-trained checkpoints: zero training cost, immediate deployment.
The primary value proposition is latency reduction for user-facing applications: MetaJuLS reduces p99 latency from 245ms to 142ms on constrained generation tasks, enabling deployment in latency-sensitive settings where baseline approaches fail SLA requirements. We provide pre-trained checkpoints at [URL] enabling deployment without retraining.
G.2 Transfer and Robustness Limitations
While rapid adaptation achieves 85–92% of specialist performance across most evaluated tasks, there are domains where transfer performs significantly worse. Specifically:
Transferring from TSP to Scheduling achieves only 74% of specialist performance
Transferring from English parsing to agglutinative languages (e.g., Turkish, Finnish) requires 15–20 gradient steps vs. 5–10 for isolating languages
Cross-domain transfer (NLP CP) shows higher variance (std 8%) compared to within-domain transfer (std 3%)
This indicates that some domain-specific knowledge may still be necessary for optimal results, particularly for domains with substantially different constraint interaction patterns.
Deterministic propagator assumption. The current approach assumes that constraint propagators are deterministic and that their effects can be accurately predicted from the current inference state. In practice, some propagators may have non-deterministic behavior or may interact in complex ways that are difficult to model. For example, approximate propagators that use randomized algorithms or propagators with emergent interaction effects from constraint composition may violate this assumption. Extending the framework to handle uncertainty in propagation effects and to reason about stochastic or approximate propagators is an important direction for future research.